Twenty-eight different video games showing violence were observed. The duration times of violence ​(in seconds) were recorded. When using this sample for a test of the claim that the population mean is greater than 85 sec, what does df​ denote, and what is its​ value?
Answer #1

As we are testing here whether the population mean is greater than 85 cc, therefore the null and the alternative hypothesis here are obtained as:

H_0: \mu = 85

H_1: \mu > 85

As we are doing the one tailed t test here to test for means, therefore the degrees of freedom for the t test here is obtained as:
Df = n - 1 = 28 - 1 = 27

Therefore 27 is the degrees of freedom here.